# Fractional binomial regression model for count data with excess zeros

### v1

@@ 3. Fractional binomial regression model
Under assumption 3.1, from Proposition 2.2 in [8] it follows that $P(B_{N}(p,H,c)=\ell)\neq 0$ for any $\ell=0,1,\cdots,N.$ Note that from (3.1-3.3) the log of the pmf of a fractional binomial distribution is continuous with respect to $(p,H,c^{\prime})$ which is, again, continuous with respect to the parameter $\Theta.$ Therefore, $\ln P(y_{i}|{\bf x}_{i},\Theta)$ is continuous in $\Theta$ for all $(y_{i},{\bf x}_{i}).$ Also, since $P(y_{i}|{\bf x}_{i},\Theta)$ is positive and continuous in $\Theta,{\bf x}_{i}$, by Assumption 3.2 i,iii, $|\ln P(y_{i}|{\bf x}_{i},\Theta)|<D(y_{i})<\infty$ for all $\Theta\in\mathcal{C},{\bf x}_{i}\in\mathcal{X}$ and $\int\sum_{y=0}^{N}P(y|{\bf x},\Theta_{0})D(y)dg(\bf{x})<\infty.$ Therefore, by Lemma 2.4 in [10]
@@ 5.1. Data from horticulture
Pho(categorical) FB VS ZINB-2 0.971 0.166
BAP(numerical) FB VS ZINB 1.226 0.110
FB VS ZIP 1.723 .042
Pho, BAP FB VS ZINB-2 1.482 .069
(categorical) FB VS ZINB .508 .306
FB VS ZIP 1.144 .126
@@ 5.2. Data from public health
logit link variable Categories Estimates (p-value) (Intercept) -1.450 (.000) Age 15-24 Ref 25-34 .211 (.043) $p$ 35-49 .046 (.759) Educational level Not educated Ref Primary .522 (.000) Secondary and higher .917 (.000) (Intercept) 11.725(.904) Age 15-24 Ref 25-34 1.368 (.992) $H$ 35-49 -1.683 (.986) Educational level Not educated Ref Primary .309 (.956) Secondary and higher -26.726 (.000) (Intercept) -1.666 (.000) Age 15-24 Ref 25-34 .054 (.805) $c^{\circ}$ 35-49 .355 (.150) Educational level Not educated Ref Primary -.460(.024) Secondary and higher -17.717 (.000)
@@
log/logit variable Categories Estimates (p-value) link ZIP ZINB ZINB-2 (Intercept) 1.136 (.000) 1.136 (.000) 1.136 (.000) Age 15-24 Ref Ref Ref $\mu$ 25-34 .116 (.086) .117(.086) .116 (.087) (log) 35-49 .129 (.156) .129(.156) .129 (.157) Educational level Not educated Ref Ref Ref Primary .123 (.064) .123(.064) .123 (.064) Secondary and higher .170 (.067) .170(.066) .171 (.067) (Intercept) -.450(.098) -.450 (.098) -.448 (.100) Age 15-24 Ref Ref Ref $\pi$ 25-34 -.205 (.481) -.205(.481) -.206 (.479) (logit) 35-49 .358 (.265) .359(.264) .358 (.267) Educational level Not educated Ref Ref Ref Primary -.954 (.000) -.954 (.000) -.955 (.000) Secondary and higher -7.261 (.814) -8.317 (.897) -6.576 (.715) (Intercept) 17.413(.075) 13.404(.930) Age 15-24 Ref $\theta$ 25-34 3.482(.969) (log) 35-49 2.462(.923) Educational level Not educated Ref Primary 2.783(.977) Secondary and higher 2.245(.969)
@@
test-statistics p-value FB VS ZINB-2 3.756 .000 FB VS ZINB 1.855 .032 FB VS ZIP 1.378 .084

### v2

@@ 3. Fractional binomial regression model
Under Assumption 3.1, it follows from Proposition 2.2 in [8] that $P(B_{N}(p,H,c)=\ell)\neq 0$ for any $\ell=0,1,\cdots,N.$ Note that from (3.1-3.3) the log of the pmf of a fractional binomial distribution is continuous with respect to $(p,H,c^{\prime})$ which is, again, continuous with respect to the parameter $\Theta.$ Therefore, $\ln P(y_{i}|{\bf x}_{i},\Theta)$ is continuous in $\Theta$ for all $(y_{i},{\bf x}_{i}).$ Also, since $P(y_{i}|{\bf x}_{i},\Theta)$ is positive and continuous in $\Theta,{\bf x}_{i}$, by Assumption 3.2 i,iii, $|\ln P(y_{i}|{\bf x}_{i},\Theta)|<D(y_{i})<\infty$ for all $\Theta\in\mathcal{C},{\bf x}_{i}\in\mathcal{X}$ and $\int\sum_{y=0}^{N}P(y|{\bf x},\Theta_{0})D(y)dg(\bf{x})<\infty.$ Therefore, by Lemma 2.4 in [10]
@@ 5.1. Data from horticulture
Pho(categorical) FB VS ZINB-2 0.97 0.17
BAP(numerical) FB VS ZINB 1.23 0.11
FB VS ZIP 1.72 0.04
Pho, BAP FB VS ZINB-2 1.48 0.07
(categorical) FB VS ZINB 0.51 0.31
FB VS ZIP 1.14 0.13
@@ 5.2. Data from public health
logit link variable Categories Estimates (p-value) (Intercept) -1.45 (0.00) Age 15-24 Ref 25-34 0.21 (0.04) $p$ 35-49 0.05 (0.76) Educational level Not educated Ref Primary 0.52 (0.00) Secondary and higher 0.92 (0.00) (Intercept) 11.73(0.90) Age 15-24 Ref 25-34 1.37 (0.99) $H$ 35-49 -1.68 (0.99) Educational level Not educated Ref Primary 0.31 (0.96) Secondary and higher -26.73 (0.00) (Intercept) -1.67 (0.00) Age 15-24 Ref 25-34 0.05 (0.81) $c^{\circ}$ 35-49 0.36 (0.15) Educational level Not educated Ref Primary -0.46(0.02) Secondary and higher -17.72 (0.00)
@@
log/logit variable Categories Estimates (p-value) link ZIP ZINB ZINB-2 (Intercept) 1.14 (0.00) 1.14 (0.00) 1.14 (0.00) Age 15-24 Ref Ref Ref $\mu$ 25-34 0.12 (0.09) 0.12(0.09) 0.12 (0.09) (log) 35-49 0.13 (0.16) 0.13(0.16) 0.13 (0.16) Educational level Not educated Ref Ref Ref Primary 0.12 (0.06) 0.12(0.06) 0.12 (0.06) Secondary and higher 0.17 (0.07) 0.17(0.07) 0.17 (0.07) (Intercept) -0.45(0.10) -0.45 (0.10) -0.45 (0.10) Age 15-24 Ref Ref Ref $\pi$ 25-34 -0.21 (0.48) -0.21(0.48) -0.21 (0.48) (logit) 35-49 0.36 (0.27) 0.36(0.26) 0.36 (0.27) Educational level Not educated Ref Ref Ref Primary -0.95 (0.00) -0.95 (0.00) -0.96 (0.00) Secondary and higher -7.26 (0.81) -8.32 (0.90) -6.58 (0.72) (Intercept) 17.41(0.08) 13.40(0.93) Age 15-24 Ref $\theta$ 25-34 3.48(0.97) (log) 35-49 2.46(0.92) Educational level Not educated Ref Primary 2.78(0.98) Secondary and higher 2.25(0.97)
@@
test-statistics p-value FB VS ZINB-2 3.76 0.00 FB VS ZINB 1.86 0.03 FB VS ZIP 1.38 0.08
@@ 6. Discussion and conclusion
Data Sets and Computer Code
R-code used in simulations and data analysis is available from GitHub repository https://github.com/leejeo25/frbinom_regression.git and https://github.com/leejeo25/fbglm.git. Dataset used in subsection 5.1 is publicly available and dataset in subsection 5.2 was obtained from the Demographic and Health Surveys (DHS) Program.